# WHALE-FL: Wireless and Heterogeneity Aware Latency Efficient Federated Learning over Mobile Devices via Adaptive Subnetwork Scheduling

### v1

@@ Abstract
As a popular distributed learning paradigm, federated learning (FL) over mobile devices fosters numerous applications, while their practical deployment is hindered by participating devices’ computing and communication heterogeneity. Some pioneering research efforts proposed to extract subnetworks from the global model, and assign as large a subnetwork as possible to the device for local training based on its full computing and communications capacity. Although such fixed size subnetwork assignment enables FL training over heterogeneous mobile devices, it is unaware of (i) the dynamic changes of devices’ communication and computing conditions and (ii) FL training progress and its dynamic requirements of local training contributions, both of which may cause very long FL training delay. Motivated by those dynamics, in this paper, we develop a wireless and heterogeneity aware latency efficient FL (WHALE-FL) approach to accelerate FL training through adaptive subnetwork scheduling. Instead of sticking to the fixed size subnetwork, WHALE-FL introduces a novel subnetwork selection utility function to capture device and FL training dynamics, and guides the mobile device to adaptively select the subnetwork size for local training based on (a) its computing and communication capacity, (b) its dynamic computing and/or communication conditions, and (c) FL training status and its corresponding requirements for local training contributions. Our evaluation shows that, compared with peer designs, WHALE-FL effectively accelerates FL training without sacrificing learning accuracy.
[…]
1 Introduction
@@
Federated Learning (FL) [20] recently experienced a notable evolution, expanding its scope from conventional data center environments to harness the potential of mobile devices [15, 18, 4]. This shift has been propelled by the continuous advancements in hardware, empowering mobile devices like the NVIDIA Xavier, Galaxy Note20, iPad Pro, MacBook laptops, etc. with increasingly robust on-device computing capabilities tailored for local training. With the collective intelligence of ending edge devices and FL’s fundamental principle of preserving data privacy, FL over mobile devices has paved the way for a diverse spectrum of innovative mobile applications, including keyboard predictions [11], smart home hazard detection [24], health event detection [3], and so on.
@@
While FL over mobile device has great potentials, its practical deployment faces significant challenges due to the inherent heterogeneity among real-world mobile devices, varying in computing capability, wireless conditions and local data distribution [14]. Existing FL studies often assume the model-homogeneous setting, where global and local models share identical architectures across all clients. However, as devices are forced to train models within their individual capability, developers have to choose between excluding low-tier devices, introducing training bias [2], or maintaining a low-complexity global model to accommodate all clients, resulting in degraded accuracy [6, 23]. The trend towards large models like Transformers [19] exacerbates the issue, hindering their training on mobile devices. Furthermore, unlike GPU clusters with stable high-speed Internet connections, mobile devices’ computing resources are constrained and heterogeneous and their wireless transmissions are relatively slow and dynamic, both of which lead to huge latency in FL training [5] and may severely degrade the performance of associated applications.
@@
To address the limitations of model-homogeneous FL, researchers have recently studied how to enable heterogeneous models across the server and mobile clients during FL training. Subnetwork training, exemplified by pioneering approaches like Federated Dropout [22], HeteroFL [9], and FjORD [13], exhibits efficacy by having mobile devices train smaller subnetworks extracted from the large global server model. These designs offer solutions to aggregating diverse devices’ subnetworks. By tailoring subnetwork architecture for the individual device, subnetwork training can ensure compatibility with mobile devices owning heterogeneous computing and communications capability. However, a prevalent challenge in current subnetwork approaches lies in their static fixed-size subnetwork assignment policy. Such a policy may fail to realize the full potential of subnetwork based training, mainly due to the unawareness of system dynamics (i.e., computing and communications dynamics) and FL training dynamics.
@@
System dynamics encompass the time-varying computing loads of devices’ background applications and the fluctuating wireless communication conditions across FL training rounds, which affects the sizes of subnetworks that a mobile device can support over rounds. Since most modern mobile devices (e.g., smartphones) participating in FL training have the ability to run multiple tasks (e.g., video streaming, image processing, and social media updates [1]) simultaneously, the dynamic orchestration of CPU/GPU resources across these concurrent activities results in the fluctuations in computing power and available memory for FL tasks, consequently impacting the supported subnetwork sizes for on-device computing. Similarly, wireless communications dynamics caused by users’ mobility, wireless channel fading, etc. lead to dynamic transmission rates, which directly affect candidate subneworks sizes that a mobile device can support for local model updates.
FL training dynamics represents FL convergence’s dynamic requirements for the contributions from local training at different training stages, which implicitly affects participating devices’ selections on subnetwork sizes. As FL training starts from scratches, any contributions from local training is helpful at the early training stage. Then, when FL training steps into the middle stage, it requires more accurate local training results to converge the global model. Finally, at the late stage, i.e., when FL training is close to the convergence, most participating devices have already made sufficient contributions to the global model and their local training may only have marginal benefits for global model’s convergence.
@@
We observe that failing to capture system or FL training dynamics and always using the possible largest-sized subnetworks under devices’ full capabilities may significantly prolong FL training process. Different from prior static fixed-size subnetwork assignment methods, in this paper, we propose a wireless and heterogeneity aware latency efficient FL (WHALE-FL) approach to accelerate FL training via adaptive subnetwork scheduling. WHALE-FL characterizes system dynamics and FL training dynamics and tailors appropriate sized subnetworks for heterogeneous mobile devices under dynamic computing/wireless environments at different FL training stages. As far as we know, WHALE-FL is the first paper that converts static fixed-size subnetwork allocation (e.g., HeteroFL [9], FjORD [13], etc.) into dynamic/adaptive subnetwork scheduling for each device by jointly considering system heterogeneity and FL training dynamics, and conducts system level experiments for validation. Our salient contributions are summarized as follows.
@@
2 Preliminary
2.1 FL over Heterogeneous Mobile Devices
Consider that $M$ mobile devices in a wireless network collaboratively engage in FL to train a deep neural network on locally distributed data sets $\{D_{1},\cdots,D_{i},\cdots,D_{M}\}$. Their local models are parameterized by $\{W_{1},\cdots,W_{i},\cdots,W_{M}\}$, which are updated using stochastic gradient descents [21] on the local data samples through local training. The server collects the local model updates and aggregates them into a global model $W_{g}$ using model averaging [20, 17]. This aggregation occurs over multiple communication rounds, with the global model at the $r$-th round denoted as $W_{g}^{r}=\frac{1}{M}\sum_{m=1}^{M}W_{m}^{r}$. In the subsequent training round, $W_{g}^{r}$ is transmitted to mobile devices, and their local models are updated as $W_{i}^{r+1}=W_{g}^{r}$. This process repeats until FL converges, while system heterogeneity (communications and computing) among mobile devices incurs huge training latency and significantly slows down FL convergence.
2.2 FL with Subnetwork Extraction
To address system heterogeneity issue in FL training, subnetwork method was introduced in [9], which extracts different sized subnetworks from global model.
Let $\mathcal{W}^{P}=\{W^{1},W^{2},\cdots,W^{p},\cdots,W^{P}\}$ be a collection of candidate subnetworks to be selected by mobile devices for local training, where $P$ complexity/size levels are considered. A lower size level $p$ corresponds to a larger-sized subnetwork, and $W^{P}$ is the smallest subnetwork for selection, i.e., $W^{P}\subset W^{P-1}\subset\cdots\subset W^{1}$. We follow the same approach as illustrated in [9] to extract subnetworks from the global model by shrinking the width of hidden channel with specific ratios. Let $s\in(0,1]$ be the hidden channel shrinkage ratio. Then, we have $|W^{p}|/|W_{g}|=|W^{p}|/|W^{1}|=s^{2(p-1)}$. With this construction, different sized subnetworks can be assigned to participating mobile devices according to their corresponding capabilities. Suppose that the number of devices in each subnetwork size level is $\{M_{1},\cdots,M_{P}\}$. The server has to aggregate the heterogeneous subnetworks in every training round. As demonstrated in [9], the global aggregation is conducted as follows.
@@ 2.2 FL with Subnetwork Extraction
In this way, each parameter is averaged from those devices whose assigned subnetwork contains that specific parameter, which enables the global aggregation and FL training with different sizes of subnetworks. Although the subnetwork method in [9] alleviates system heterogeneity issue, it is a fixed policy. It cannot capture the dynamic changes of wireless transmission/on-device computing conditions, or the dynamic requirements of contributions from local training at different FL training stages, either of which may result in a huge FL training latency.
3 Motivation
@@
Unawareness of system dynamics. Traditional subnetwork assignment (e.g., HeteroFL [9]) is fixed, which is based on the participating mobile device’s maximum system capability (i.e., computing + communications), while ignoring the dynamic changes of the device’s computing and communication conditions. Such an unawareness may lead to poor subnetwork assignment decisions and significantly delay the FL training process. For instance, a mobile device capable of computing a full-sized model may be experiencing a bad wireless access (e.g., 4G/LTE) or running some computing intensive background applications (e.g., GPU intensive gaming) in a certain training round. In this case, the fixed full-sized subnetwork assignment may become this device a straggler and cause a big latency in FL training. Thus, an adaptive subnetwork scheduling aware of system (computing + communication) dynamics is in need.
@@
Unawareness of FL training dynamics. The fixed subnetwork assignment is unaware of FL training progress and its dynamic requirements of learning contributions from local mobile devices. Specifically, in early FL training stage, FL training starts from scratches and any contributions from any device’s local training will be helpful, regardless of the subnetwork size. At this stage, using small-sized subnetworks can expedite on-device computing and wireless transmissions of local model updates, which helps to reduce the FL training latency. Then, as FL training proceeds, say the middle stage, more and more accurate local model updates are needed for global training model to converge. Thus, only using small-sized subnetwork for speed while sacrificing local learning performance is not a wise option for mobile devices, since it may slow down or even fail FL convergence. Further, when FL training is close to the convergence (i.e., the late stage), most mobile devices have already made substantial/enough contributions to the global model. As for those devices, sticking to the large/full-sized subnetwork for local training has very limited learning benefits for FL convergence, while some computing/communications constrained ones among those devices may incur huge training latency or even become stragglers. Therefore, it is necessary to develop an adaptive subnetwork scheduling that captures FL training dynamics, recognizes computing/communication constraints and selects right sized subnetworks for local training, to improve the delay efficiency in FL training over mobile devices.
4 WHALE-FL Design
@@
Figure 1: The conceptual sketch of WHALE-FL.
4.1 Adaptive Subnetwork Selection Utility
WHALE-FL’s adaptive subnetwork selection performance hinges on two critical aspects: system efficiency and training efficiency. System efficiency encompasses the duration of each training round, including local computing and model updates time consumption. Training efficiency gauges the local training’s contributions to global convergence. The fluctuating wireless conditions and available computing resources of devices, as well as their training progress with local data, collectively determine the system and training efficiency, forming what we term as adaptive subnetwork selection utility.
To accelerate FL training without sacrificing learning accuracy, it is critical to trade-off system and training efficiencies to select the appropriate subnetwork size for individual device’s local training per round. Briefly, WHALE-FL favors system efficiency over training efficiency at the early stage of FL training, and tends to schedule small-sized subnetworks for devices’ local training. While FL training steps into the middle stage, if more accurate local training is needed for FL convergence, WHALE-FL prefers training efficiency to system efficiency and schedules to adaptively increase the size of subnetworks for participating mobile devices. Otherwise, WHALE-FL prioritizes system efficiency over training efficiency. When FL is close to convergence, WHALE-FL jointly considers system and training efficiencies, and gradually decreases the size of subnetworks for local training, given the fact that most devices have contributed enough to global model and it is unnecessary to keep large-sized subnetworks for local training.
@@ 4.1 Adaptive Subnetwork Selection Utility
where $T_{i,r}^{tr}$ and $T_{i,r}^{co}$ are the transmission delay and the computing delay, respectively, for the unit/smallest subnetwork. We assume that the wireless transmission rates and available computing resources dynamically change over rounds, but are relatively stable within a FL training round. Thus, given a learning task, transmission and computing workloads for the unit subnetwork are fixed, and $T_{i,r}^{tr}$ and $T_{i,r}^{co}$ can be easily estimated for device $i$ in the $r$-th round. A higher $SE_{i,r}$ enables devices to opt for larger subnetwork sizes for local training within this round, and vice versa. The formulation in Eqn. (3) comprehensively covers the system efficiency for communication delay dominant cases (i.e., slow transmissions & fast computing), computing delay dominant cases (i.e., fast transmissions & slow computing), and communication-computing comparable cases.
Training efficiency utility. WHALE-FL employs training efficiency utility to characterize FL training dynamics. WHALE-FL leverages training loss that measures the estimation error between model predictions and the ground truth to identify the current training status. Let $l_{i,r}$ denote the average training loss on the samples over all the local iterations in the $r$-th round, which can be automatically generated during on-device training with negligible collection overhead. In an adaptive update manner, we define the training efficiency $TE_{i,r}$ for device $i$ in the $r$-th round as follows:
$TE_{i,r}=\quad TE_{i,r-1}\quad+$
$\gamma\cdot[2\cdot\mathbbm{1}(l_{i,r-1}\geq\alpha\cdot l_{target})-1]\cdot\mathbbm{1}(\Delta l_{i,r-1}\leq\Delta_{th}),$ (4)
where $l_{target}$ denotes the target loss depending on the specific FL training task, $\mathbbm{1}(x)$ is an indicator function that takes value $1$ if the statement $x$ is true and 00 otherwise, $\alpha\geq 1$ is a developer-defined constant governing the delineation of training stages, and $\gamma\geq 0$ is a hyper-parameter controlling the step-size of utility increase or decrease compared to the previous round. Besides, $\Delta l_{i,r}=l_{i,r}-l_{i,r-1}$ denotes the gradient of local training loss at device $i$ in the $r$-th round, and $\Delta_{th}$ is the developer-specified threshold for local training loss gradient. Here, $\mathbbm{1}(l_{i,r-1}\geq\alpha\cdot l_{target})$ represents whether the FL training is close to the convergence, and $\mathbbm{1}(\Delta l_{i,r-1}\leq\Delta_{th})$ represents whether device $i$’s local training can make good statistical contributions [14] to FL training.
At the early FL training stage, $(l_{i,r-1}\geq\alpha\cdot l_{target})$ and $(\Delta l_{i,r-1}>\Delta_{th})$, which indicates that device $i$ makes sufficient contributions to FL training, and should keep the subnetwork size from training efficiency perspective. Thus, following Eqn. (4.1), $TE_{i,r}=TE_{i,r-1}$. At the middle FL training stage, there are two cases: Case I: $(l_{i,r-1}\geq\alpha\cdot l_{target})$ and $(\Delta l_{i,r-1}\leq\Delta_{th})$, which means device $i$ cannot make good enough contributions to FL training with current size of subnetwork, and should adaptively increase the subnetwork size from training efficiency perspective. Thus, following Eqn. (4.1), $TE_{i,r}=TE_{i,r-1}+\gamma$; Case II: $(l_{i,r-1}\geq\alpha\cdot l_{target})$ and $(\Delta l_{i,r-1}>\Delta_{th})$, which means after increasing the subnetwork size in Case I, device $i$ can make good contributions to FL training again, and should keep the subnetwork size from training efficiency perspective. Thus, following Eqn. (4.1), $TE_{i,r}=TE_{i,r-1}$. Device $i$ may switch between those two cases at the middle FL training stage. At the late FL training stage when FL is about to converge, $(l_{i,r-1}<\alpha\cdot l_{target})$, i.e., $l_{i,r-1}$ stabilizes at a small value, and $(\Delta l_{i,r-1}\leq\Delta_{th})$. That indicates device $i$ has already made substantial contributions to FL training and sticking to the large size subnetwork for $i$’s local training has very limited benefits for FL convergence. Thus, device $i$ can gradually downsize its local subnetworks from training efficiency perspective, i.e., $TE_{i,r}=TE_{i,r-1}-\gamma$. As such, Eqn. (4.1) can characterize FL training dynamics and adjust the training efficiency utility values for devices’ adaptive subnetwork selection at different training stages.
$\begin{split}Util(i,r)=\underbrace{\left({T_{i,r}^{tr}+T_{i,r}^{co}}\right)^{-1}}_{\textit{System efficiency utility}}\times\,\,\,{\underbrace{\left\{TE_{i,r-1}+\gamma\cdot\left[2\cdot\mathbbm{1}(l_{i,r-1}\geq\alpha\cdot l_{target})-1\right]\cdot\mathbbm{1}(\Delta l_{i,r-1}\leq\Delta_{th})\right\}}_{\textit{Training efficiency utility}}}^{\beta}\end{split}$ (5)
@@
Adaptive subnetwork selection utility function. WHALE-FL trades-off the system and training efficiencies to determine the utility values for subnetwork scheduling over rounds. The adaptive subnetwork selection utility function is shown in Eqn. (5), where $Util(i,r)$ associates system and training efficiencies with developer-specified factor $\beta$. Aware of both system and FL training dynamics, a large/small value of $Util(i,r)$ suggests that device $i$ should opt for a large/small sized subnetwork in the subsequent $r$-th round.
4.2 Utility Value to Subnetwork Size Conversion
@@
The calculated utility in Eqn. (5) cannot directly be used by individual mobile device to decide its subnetwork size selection. To facilitate mobile devices’ decisions, it is necessary to convert subnetwork selection utility values into available/candidate subnetwork sizes.
@@
Given definitions above, the next step is to normalize devices’ utility values into the range of $[0,1]$, in order to identify the model shrinkage ratio. We propose to use a piecewise linear function to normalize $Util(i,r)$ into $U_{n}(i,r)$ as follows.
@@
Here, $|W^{max}_{i}|$ denotes the maximum subnetwork size that device $i$ can support with its full computing capacity, where $W^{max}_{i}\in\mathcal{W}^{P}$ as defined in Sec. 2.2. $\hat{W}(i,r)\in\mathcal{W}^{P}$ is a subnetwork derived from normalized utility value $U_{n}(i,r)$, which can be expressed as

### v3

@@ Abstract
As a popular distributed learning paradigm, federated learning (FL) over mobile devices fosters numerous applications, while their practical deployment is hindered by participating devices’ computing and communication heterogeneity. Some pioneering research efforts proposed to extract subnetworks from the global model, and assign as large a subnetwork as possible to the device for local training based on its full computing capacity. Although such fixed size subnetwork assignment enables FL training over heterogeneous mobile devices, it is unaware of (i) the dynamic changes of devices’ communication and computing conditions and (ii) FL training progress and its dynamic requirements of local training contributions, both of which may cause very long FL training delay. Motivated by those dynamics, in this paper, we develop a wireless and heterogeneity aware latency efficient FL (WHALE-FL) approach to accelerate FL training through adaptive subnetwork scheduling. Instead of sticking to the fixed size subnetwork, WHALE-FL introduces a novel subnetwork selection utility function to capture device and FL training dynamics, and guides the mobile device to adaptively select the subnetwork size for local training based on (a) its computing and communication capacity, (b) its dynamic computing and/or communication conditions, and (c) FL training status and its corresponding requirements for local training contributions. Our evaluation shows that, compared with peer designs, WHALE-FL effectively accelerates FL training without sacrificing learning accuracy.
Introduction
@@
Federated Learning (FL) (McMahan et al. 2017) recently experienced a notable evolution, expanding its scope from conventional data center environments to harness the potential of mobile devices (Li et al. 2021a; Chen et al. 2023). This shift has been propelled by the continuous advancements in hardware, empowering mobile devices like the NVIDIA Xavier, iPhone 16, etc. with increasingly robust on-device computing capabilities tailored for local training. With the collective intelligence of edge devices and FL’s fundamental principle of preserving data privacy, FL over mobile devices has paved the way for a diverse spectrum of innovative mobile applications, including keyboard predictions (Hard et al. 2018), smart home hazard detection (Yu et al. 2020), health event detection (Brisimi et al. 2018), and so on.
@@
While FL over mobile device has great potentials, its practical deployment faces significant challenges due to the inherent heterogeneity among real-world mobile devices, varying in computing capability, wireless conditions and local data distribution (Lai et al. 2021). Existing FL studies often assume the model-homogeneous setting, where global and local models share identical architectures across all clients. However, as devices are forced to train models within their individual capability, developers have to choose between excluding low-tier devices, introducing training bias (Bickel, Hammel, and O’Connell 1975), or maintaining a low-complexity global model to accommodate all clients, resulting in degraded accuracy (Cho, Wang, and Joshi 2021; Ye et al. 2020). The trend towards large models like Transformers (Liu et al. 2023) exacerbates the issue, hindering their training on mobile devices. Furthermore, unlike GPU clusters with stable high-speed Internet connections, mobile devices’ computing resources are constrained and heterogeneous, and their wireless transmissions are relatively slow and dynamic, both of which lead to huge latency in FL training (Chen et al. 2022) and may severely degrade the performance of associated applications.
@@
To address the limitations of model-homogeneous FL, researchers have recently studied how to train different sized models across heterogeneous mobile clients and corresponding global model aggregation in FL training. Subnetwork training, exemplified by pioneering approaches like width-based subnetwork generation in Federated Dropout (Wen, Jeon, and Huang 2021) and HeteroFL (Diao, Ding, and Tarokh 2021), and depth-based generation in DepthFL (Kim et al. 2023), has proven effective by enabling mobile devices to train smaller subnetworks derived from the large global server model. These designs also offer solutions to aggregating diverse devices’ subnetworks. By tailoring subnetwork architecture for the individual device, subnetwork training can ensure compatibility with mobile devices owning heterogeneous computing and communication capabilities. However, a prevalent challenge in current subnetwork approaches lies in their static fixed-size subnetwork assignment policy. Such a policy may fail to unleash the full potential of subnetwork based training, mainly due to the unawareness of system dynamics (i.e., computing and communications dynamics) and FL training dynamics.
@@
System dynamics encompass the time-varying computing loads of devices’ background applications and the fluctuating wireless communication conditions across FL training rounds, which affects the sizes of subnetworks that a mobile device can support over rounds. Since most modern mobile devices (e.g., smartphones) participating in FL training have the ability to run multiple tasks simultaneously (Banabilah et al. 2022), the dynamic orchestration of CPU/GPU resources across these concurrent activities results in the fluctuations in computing power and available memory for FL tasks, consequently impacting the supported subnetwork sizes for on-device computing. Similarly, wireless communications dynamics caused by users’ mobility, wireless channel fading, etc. lead to dynamic transmission rates, which directly affect candidate subnework sizes that a mobile device can support for local model updates.
FL training dynamics represents FL convergence’s dynamic requirements for the contributions from local training at different training stages, which implicitly affects participating devices’ selections on subnetwork sizes. Recent studies have revealed that critical learning periods (CLP) exist in the training process of deep neural networks (Achille, Rovere, and Soatto 2018; Yan, Wang, and Li 2022). As the FL training proceeds, the training contributions from each client gradually decrease. Thus, it is crucial to assign suitable subnetwork sizes to clients based on FL training dynamics.
@@
We observe that failing to capture system or training dynamics and always using the possible largest-sized subnetworks under devices’ full capabilities may significantly prolong the FL training process. Different from prior static fixed-size subnetwork assignment methods, in this paper, we propose a wireless and heterogeneity aware latency efficient FL (WHALE-FL) approach to accelerate FL training via adaptive width-wise subnetwork scheduling. WHALE-FL characterizes system dynamics and FL training dynamics and tailors appropriate-sized subnetworks for heterogeneous mobile devices under dynamic computing/wireless environments at different FL training stages. As far as we know, WHALE-FL is the first paper that converts static fixed-size subnetwork allocation, e.g., HeteroFL (Diao, Ding, and Tarokh 2021), Federated Dropout (Wen, Jeon, and Huang 2021), etc., into adaptive subnetwork scheduling for each device by jointly considering system heterogeneity and FL training dynamics, and conducts system-level experiments for validation. Our salient contributions are summarized as follows:
@@
Preliminary
FL over Heterogeneous Mobile Devices
Consider that $M$ mobile devices in a wireless network collaboratively engage in FL to train a deep neural network on locally distributed datasets $\{L_{1},\cdots,L_{i},\cdots,L_{M}\}$. Their local models are parameterized by $\{W_{1},\cdots,W_{i},\cdots,W_{M}\}$, which are updated using stochastic gradient descents (Ruder 2016) on the local data samples through local training. The server collects the local model updates and aggregates them into a global model $W_{g}$ using model averaging (McMahan et al. 2017; Li et al. 2020). This aggregation occurs over multiple communication rounds, with the global model at the $r$-th round denoted as $W_{g}^{r}=\frac{1}{M}\sum_{m=1}^{M}W_{m}^{r}$. In the subsequent training round, $W_{g}^{r}$ is transmitted to mobile devices, and their local models are updated as $W_{i}^{r+1}=W_{g}^{r}$. This process repeats until FL converges, while system heterogeneity (communications and computing) among mobile devices incurs huge training latency and significantly slows down FL convergence.
FL with Subnetwork Extraction
To address the system heterogeneity issue in FL training, the subnetwork training method was introduced in (Diao, Ding, and Tarokh 2021), which extracts subnetworks in different sizes from the global model.
Let $\mathcal{W}^{P}=\{W^{1},W^{2},\cdots,W^{p},\cdots,W^{P}\}$ be a collection of candidate subnetworks to be selected by mobile devices for local training, where $P$ complexity/size levels are considered. A lower size level $p$ corresponds to a larger-sized subnetwork, and $W^{P}$ is the smallest subnetwork for selection, i.e., $W^{P}\subset W^{P-1}\subset\cdots\subset W^{1}$. We follow the same approach as illustrated in (Diao, Ding, and Tarokh 2021) to extract subnetworks from the global model by shrinking the width of hidden channel with specific ratios. Let $s\in(0,1]$ be the hidden channel shrinkage ratio. Then, we have $|W^{p}|/|W_{g}|=|W^{p}|/|W^{1}|=s^{2(p-1)}$. With this construction, different sized subnetworks can be assigned to participating mobile devices according to their corresponding capabilities. Suppose that the number of devices in each subnetwork size level is $\{M_{1},\cdots,M_{P}\}$. The server has to aggregate the heterogeneous subnetworks in every training round. As demonstrated in (Diao, Ding, and Tarokh 2021), the global aggregation is conducted as follows.
@@
In this way, each parameter is averaged from those devices whose assigned subnetwork contains that specific parameter, which enables the global aggregation and FL training with different sizes of subnetworks. Although the subnetwork method in (Diao, Ding, and Tarokh 2021) alleviates the system heterogeneity issue, it is a fixed policy. It cannot capture the dynamic changes of wireless transmission/on-device computing conditions, or the dynamic requirements of contributions from local training at different FL training stages, either of which may result in a huge training latency.
Fisher Information
Fisher information is utilized as a measurement of how much a change in weights can affect the output of neural networks (Achille, Rovere, and Soatto 2018). Fisher information is a 2nd-order approximation of the Hessian of the loss function (Amari and Nagaoka 2000; Martens 2014), providing information on the curvature of the loss landscape near the current weights. Such characteristics help to indicate how fast the gradient changes during training, which may be used to characterize the training dynamics from local device side and further help clients decide how to adjust their subnetwork sizes.
To enable distributed subnetwork scheduling, we use the Federated Fisher Information Matrix (FedFIM) from  (Yan, Wang, and Li 2022) instead of the traditional definition of the Fisher Information Matrix (FIM) for centralized training to avoid requiring access to the entire dataset. That is, given that training data resides in each client, the local FIM on client $i$ in the $r$-th training round is calculated by
$FI_{i,r}=\mathbb{E}_{x_{i}\sim\mathcal{X}_{i}}\mathbb{E}_{\hat{y}\sim p_{W}(\hat{y_{i}}|x_{i})}[\bigtriangledown(x_{i},\hat{y}_{i})\bigtriangledown(x_{i},\hat{y}_{i})^{\intercal}],$ (3)
where $x_{i}$ is the input data of and $y_{i}$ is the corresponding output label of client $i$, $W$ is the weight and $p_{W}(\hat{y}_{i}|x_{i})$ is the approximate posterior distribution. $\mathcal{X}_{i}$ is the empirical distribution of the $i$-th client’s local data. The corresponding gradient of the loss for $(x,y)$ is denoted as $\bigtriangledown(x,y)=\frac{\partial}{\partial W}\ell(x,y;W)$, and $\hat{y_{i}}$ is a random variable rather than a true label with its distribution following $p_{W}(\hat{y}_{i}|x_{i})$.
Motivation
@@
Unawareness of system dynamics. Traditional subnetwork assignment (e.g., HeteroFL (Diao, Ding, and Tarokh 2021)) is fixed, which is based on the participating mobile device’s maximum system capability (i.e., computing + communications), while ignoring the dynamic changes of the device’s computing and communication conditions. Such an unawareness may lead to poor subnetwork assignment decisions and significantly delay the FL training process. For instance, a mobile device capable of computing a full-sized model may be experiencing poor wireless access (e.g., 4G/LTE) or running some computing-intensive background applications (e.g., GPU-intensive gaming) in a certain training round. In this case, the fixed full-sized subnetwork assignment may make this device a straggler and cause a big latency in FL training. Thus, an adaptive subnetwork scheduling aware of system (computing + communication) dynamics is in need.
@@
Unawareness of FL training dynamics. The fixed subnetwork assignment is unaware of FL training progress and its dynamic requirements of learning contributions from local mobile devices. Since FL training starts from scratch, any contributions from any device’s local training will be helpful. Using small-sized subnetworks can expedite on-device computing and wireless transmissions of local model updates. As FL training proceeds into the CLP, more accurate local model updates are needed for the global training model to converge. When FL training is close to convergence (i.e., the late stage), most mobile devices have already made substantial contributions to the global model. For those devices, sticking to large or full-sized subnetworks for local training offers limited learning benefits for FL convergence, while some computing/communications-constrained devices may incur significant training latency or even become stragglers. Therefore, it is necessary to develop an adaptive subnetwork scheduling method that captures FL training dynamics, recognizes computing/communication constraints, and selects appropriately sized subnetworks for local training, to improve delay efficiency in FL training over mobile devices.
WHALE-FL Design
@@
Adaptive Subnetwork Selection Utility
WHALE-FL’s adaptive subnetwork selection performance hinges on two critical aspects: system efficiency and training efficiency. System efficiency encompasses the duration of each training round, including local computing and model uploading time consumption. Training efficiency gauges the local training’s contributions to global convergence. The fluctuating wireless conditions and available computing resources of devices, as well as their training progress with local data, collectively determine the system and training efficiency, forming what we term as adaptive subnetwork selection utility.
To accelerate FL training without sacrificing learning accuracy, it is critical to trade-off system and training efficiencies to select the appropriate subnetwork size for individual device’s local training per round. Briefly, WHALE-FL favors system efficiency over training efficiency at the early stage of FL training, and tends to schedule small-sized subnetworks for devices’ local training. While FL training steps into the middle stage, if more accurate local training is needed for FL convergence, WHALE-FL prefers training efficiency to system efficiency and schedules to adaptively increase the size of subnetworks for participating mobile devices. Otherwise, WHALE-FL prioritizes system efficiency over training efficiency. When FL is close to convergence, WHALE-FL jointly considers system and training efficiencies, and gradually decreases the size of subnetworks for local training, given the fact that most devices have contributed enough to the global model and it is unnecessary to keep large-sized subnetworks for local training.
@@
where $T_{i,r}^{tr}$ and $T_{i,r}^{co}$ are the transmission delay and the computing delay, respectively, for the unit/smallest subnetwork. $T$ is the developer-preferred duration of each round, which may vary for different FL tasks. We assume that the wireless transmission rates and available computing resources dynamically change over rounds, but are relatively stable within a FL training round. Thus, given a learning task, transmission and computing workloads for the unit subnetwork are fixed, and $T_{i,r}^{tr}$ and $T_{i,r}^{co}$ can be easily estimated for device $i$ in the $r$-th round. A higher $SE_{i,r}$ enables devices to opt for larger subnetwork sizes for local training within this round, and vice versa. The formulation in Eqn. (4) comprehensively covers the system efficiency for communication delay dominant cases (i.e., slow transmissions & fast computing), computing delay dominant cases (i.e., fast transmissions & slow computing), and communication-computing comparable cases.
Training efficiency utility. By employing fisher information $FI$, we define the training efficiency utility $TE_{i,r}$ for device $i$ in the $r$-th round as follows:
$TE_{i,r}=|\mathbf{B}_{i}|\sqrt{\frac{1}{|\mathbf{B}_{i}|}\sum_{k\in\mathbf{B}_{i}}\sum\limits_{d\in\mathbf{D}}\frac{FI_{i,r-d}(k)^{2}}{|\mathbf{D}|}},$ (5)
where $B_{i}$ represents the batched datasets for device $i$. Here, we utilize a window-averaged local Fisher information to measure the dynamic utility during training with $\mathbf{D}=\{1,..,d,..,D\}$ as the set of window sizes. Here, the sliding window operation helps to prevent frequent zigzag changes in subnetwork sizes, as Fisher information across different local training iterations within the $i$-th round may be highly unstable, and directly using the Fisher information of each iteration could result in unstable subnetwork selection strategies (Achille, Rovere, and Soatto 2018).
@@
Adaptive subnetwork selection utility function. WHALE-FL trades-off the system and training efficiencies to determine the utility values for subnetwork scheduling over rounds. The adaptive subnetwork selection utility function is shown in Eqn. (Adaptive Subnetwork Selection Utility), where $Util(i,r)$ associates system and training efficiencies with developer-specified factor $\beta$. Aware of both system and FL training dynamics, a large/small value of $Util(i,r)$ suggests that device $i$ should opt for a large/small sized subnetwork in the subsequent $r$-th round.
$Util(i,r)=$
${\underbrace{|\mathbf{B}_{i}|\sqrt{\frac{1}{|\mathbf{B}_{i}|}\sum_{k\in\mathbf{B}_{i}}\sum\limits_{d\in\mathbf{D}}\frac{FI_{i,r-d}(k)^{2}}{|\mathbf{D}|}}}_{\textit{Training efficiency utility}}}\,\times\underbrace{\left(\frac{T}{T_{i,r}^{tr}+T_{i,r}^{co}}\right)^{\beta}}_{\textit{System efficiency utility}}.$ (6)
Utility Value to Subnetwork Size Conversion
@@
The calculated utility in Eqn. (Adaptive Subnetwork Selection Utility) cannot directly be used by individual mobile devices to decide their subnetwork size selection. To facilitate mobile devices’ decisions, it is necessary to convert subnetwork selection utility values into available/candidate subnetwork sizes.
@@
Given the definitions above, the next step is to normalize devices’ utility values into the range of $[0,1]$, in order to identify the model shrinkage ratio. We propose to use a piecewise linear function to normalize $Util(i,r)$ into $U_{n}(i,r)$ as follows.
@@
Here, $|W^{max}_{i}|$ denotes the maximum subnetwork size that device $i$ can support with its full computing capacity, where $W^{max}_{i}\in\mathcal{W}^{P}$ as defined in Sec. FL with Subnetwork Extraction. $\hat{W}(i,r)\in\mathcal{W}^{P}$ is a subnetwork derived from normalized utility value $U_{n}(i,r)$, which can be expressed as